OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8). A. y = -2x - 4 B. y = -2x - 8 C. y = -2x + 18 D. y = -2x – 12

OpenStudy (anonymous):

use y=mx+b :) we know that m= slope, which they give you. so slope=m=-2 :) and now you need to find the y-intercept. aka "b" so plug in the values they give you like this :) -8=(-2)(2) + b so you get -8 = -4 + b isolate the b by adding 4 to both sides :) and what do you get? :) from there, use your "b" value and add it into the equation! Your answer will be y=-2x _________ we don't know if it will be + b or -b yet! solve and you will know! :D